# Multimodal Infusion Tuning for Large Models

### v1

@@ Abstract
Recent advancements in large-scale models have showcased remarkable generalization capabilities in various tasks. However, integrating multimodal processing into these models presents a significant challenge, as it often comes with a high computational burden. To address this challenge, we introduce a new parameter-efficient multimodal tuning strategy for large models in this paper, referred to as Multimodal Infusion Tuning (MiT). MiT leverages decoupled self-attention mechanisms within large language models to effectively integrate information from diverse modalities such as images and acoustics. In MiT, we also design a novel adaptive rescaling strategy at the head level, which optimizes the representation of infused multimodal features. Notably, all foundation models are kept frozen during the tuning process to reduce the computational burden(only 2.5% parameters are tunable). We conduct experiments across a range of multimodal tasks, including image-related tasks like referring segmentation and non-image tasks such as sentiment analysis. Our results showcase that MiT achieves state-of-the-art performance in multimodal understanding while significantly reducing computational overhead(10% of previous methods). Moreover, our tuned model exhibits robust reasoning abilities even in complex scenarios.
Figure 1: The overview of our proposed MiT for different modality data and tasks. For image-involved data, a pretrained CLIP encoder is employed to extract the image representation(left). Otherwise, a trainable network is employed to generate the modality embedding(such as acoustic and facial features, right). The multimodal embedding is infused into the pretrained LLM by MiT module. All of the pretrained models are kept frozen during the whole procedure.
@@ 1 Introduction
Leveraging the vast scale of model sizes and training corpus, large language models (LLMs) have showcased formidable capabilities across various real-world tasks. The prominence of LLMs is evident in the widespread attention they have garnered from both academia and industry, with notable examples being OPT [37] and LLaMA [25]. However, training a new large model is highly resource-intensive, leading to the development of numerous parameter-efficient fine-tuning (PEFT) methods, such as LLaMA-adapter [36] and (IA)${}^{3}$ [19].
@@
Currently, the majority of PEFT approaches are tailored for text-only data [15, 19], posing a challenge in handling multimodal information. Empowering LLMs with multimodal data, such as images and acoustics, remains a significant challenge. To this end, some researchers have proposed integrating image information into LLMs through PEFT, exemplified by approaches like FROMAGe [12] and mPLUG-Owl [31]. Most of these methods tend to concatenate or prefix the vision embedding with text tokens and feed them directly to the LLMs. However, we identify two drawbacks in this approach. Firstly, the concatenation of tokens significantly increases memory consumption, growing quadratically due to the self-attention mechanism’s $L^{2}$ memory requirement (where $L$ represents the length of tokens). Secondly, direct prefix approaches are deemed coarse-grained, leading to insufficient interactions among representations from multiple modalities.
@@
To address these issues, we propose a new tuning framework termed multimodal infusion tuning (MiT). Specifically, we decouple the self-attention of large language models and progressively infuse the global representations of other modalities into the keys ($K$) and values ($V$) of text embedding. The infusion procedure is designed to be linear, ensuring low memory consumption. Additionally, we introduce a novel adaptive rescaling strategy at the head level to enhance collaboration between different modalities. This infusion strategy allows us to endow LLMs with the capability to process multimodal information. One major difference between MiT and other methods is that we decouple self-attention in-depth, and inject multimodal information at a more fine-grained aspect. We freeze the parameters in pretrained models and exclusively fine-tune those newly introduced in MiT(around only 2.5% of all parameters). Unlike previous approaches, MiT is versatile and can handle not only image data but also other types of data, such as acoustic and facial features. For various downstream tasks, we employ different head networks during tuning, such as a lightweight decoder for referential segmentation and a fully connected layer for sentiment analysis. We conduct experiments on seven datasets encompassing three tasks, and the results demonstrate that MiT achieves state-of-the-art performance while maintaining efficiency(our proposed method requires only 0.47 TFLOPs).
In summary, our paper presents three key contributions:
@@
  • We introduce a fine-grained tuning strategy for LLMs, named multimodal infusion tuning. This strategy is designed to efficiently accommodate various types of multimodal data, allowing for their progressive infusion into LLMs.
@@ 2 Related Works
2.1 Large Language Models
Large language models (LLMs) have demonstrated significant abilities in in-context learning and long-term generation recently, exemplified by models such as ChatGPT, OPT [37], and LLaMA [25]. To enhance the application of LLMs in downstream tasks, various methods have been introduced, including adapter-tuning [7] and prefix-tuning [15]. These methods primarily focus on fine-tuning the model while preserving the original parameters, yielding two key benefits: reduction in computational overhead and preservation of the LLM’s proficiency trained on large corpora. More recently, novel Prompt Encoding for Fine-Tuning (PEFT) methods have emerged, such as LLaMA-adapter [36] and (IA)${}^{3}$ [19]. Notably, (IA)${}^{3}$ introduces learnable vectors that are multiplied with the keys (K) and values (V) in self-attention mechanisms, thereby reducing memory usage. Our proposed method draws considerable inspiration from (IA)${}^{3}$, albeit specifically tailored for multimodal data and tasks.
@@ 3 Multimodal Infusion Tuning
Our method aims to seamlessly integrate multimodal information into a pretrained LLM while keeping its parameters frozen. The overall pipeline is illustrated in Figure 1. We introduce a tunable MiT module into the pretrained LLM, facilitating the integration of representations from other modalities. Through tuning, the LLM progressively acquires the capability to process multimodal signals as MiT module is zero-initialized. The resulting interacted textual features, combined with visual features in tasks involving images(Figure 1-left), are then fed into various downstream networks for predictions. Consequently, our method finds applications across diverse domains, including referent segmentation, image-text classification, and sentiment analysis.
Figure 2: The detailed architecture of MiT. We decouple the self-attention and integrate the image representation LLM via element-wise multiplication and addition. The integration is performed on both self-attention(left) and feed-forward(right) module.
@@ 3.1 Architecture Design
As illustrated in Figure 2, we infuse the multimodal information in both self-attention and feed-forward module of pretrained LLMs. We design MiT as a plugable module, so we can leverage the capabilities of the LLM learnt from large scale corpus.
Without loss of generality, assume that we have an global image representation extracted by CLIP image encoder [22]: $I\in R^{d_{I}}$, where $d_{I}$ is the corresponding feature dimension. We delve into textual self-attention and integrate $I$ into key($K_{t}$) and value($V_{t}$). Take $V_{t}$ for example, to fit it into the space of text representations, we employ two transformations:
@@
where $W^{v}_{d},W^{v}_{a}\in R^{d_{I}\times d_{T}}$, $b^{v}_{d},b^{v}_{a}\in R^{d_{T}}$ are learnable vectors. Then we infuse the transformed representation into $V_{t}$ via element-wise multiplication and addition:
@@
where $I_{d}^{k}$ and $I_{a}^{k}$ are generated in the same approach as Equation 1.
@@ 3.2 Downstream Task Tuning
We fine-tune our MiT on three tasks for multimodal understanding: refer segmentation, image-text classification, and multimodal sentiment analysis. Among these tasks, sentiment analysis does not involve images, while the other two require image data. For tasks involving images, we utilize a frozen CLIP encoder to generate global semantic representations and features of images at different levels, as illustrated in Figure 1. Alternatively, for tasks not involving images, we employ a trainable transformer feature encoder to extract multimodal embeddings. Since the LLMs employed in our approach are decoder-only structures, we extract the semantic representations for downstream tasks from the last token of the entire sequences.
For refer segmentation, the goal is to segment the object from the image according to a descriptive. Under this scenario, we design a light-weight decoder, which takes infused text embedding and multi-level image features as input, and generate the predicted mask. The DICE loss [21] is employed in our framework:
@@ 3.3 Implementation Details
We evaluate our method on RefCOCO [32], RefCOCO+ [32], and G-Ref [20] for refer segmentation. There are 19994, 19992, and 26711 images, with 50000, 49856, and 54822 references, respectively. For image-text classification, we employ two datasets: UPMC-Food101 [27] and SNLI-VE [29], which contain 90840 and 565286 image-text pairs, respectively. UPMC-Food101 provides 101-category classification task while SNLI-VE aims to reason the semantic relationship(entailment, neutral, or contradiction) between text hypothesis and image premise. As for the sentiment analysis, we use two popular dataset: MOSI [33] and MOSEI [35], in which the sentiment tendency are annotated within [-3, +3]. The goal is to predict the sentiment tendency based on corresponding texts, acoustic and facial features.
We use the popular LLaMA [25] with 7B parameters as the textual foundation model. For image branch, we employ the CLIP ViT-L-336 model [22] to generate the global representations. In MiT, we infuse the multimodal representation into last certain layers in LLM, and the detailed effect are discussed in Section 5.1. There are three layers in the light-weight decoder for refer segmentation to maintain efficiency. In sentiment analysis task, three-layer transformer blocks are employed to encoder acoustic and facial features. Inherited from pretrained models, $d_{I}$, $d_{T}$, and $d_{T^{\prime}}$ are set to 768, 4096, and 11008, respectively.
@@ 4.1 Results on Referring Segmentation
Table 2: This results of our method on image-text classification benchmarks. The accuracy is utilized as metrics in the table.
@@ 5.1 Ablation Studies
Modules of MiT. There are several components in our proposed MiT, including K/V infusion, adaptive rescaling, feed-forward infusion, etc. The ablation study of each part on RefCOCO benchmark are shown in Table 4. As we can see from the results, K/V infusion plays the most important role in MiT. When the adaptive head rescaling is removed, the performance of the model will also decrease accordingly, demonstrating the effect of this part. Besides, infusing in feed-forward module also plays a positive role for the final results.

### v2

@@ Abstract
Recent advancements in large-scale models have showcased remarkable generalization capabilities in various tasks. However, integrating multimodal processing into these models presents a significant challenge, as it often comes with a high computational burden. To address this challenge, we introduce a new parameter-efficient multimodal tuning strategy for large models in this paper, referred to as Multimodal Infusion Tuning (MiT). MiT leverages decoupled self-attention mechanisms within large language models to effectively integrate information from diverse modalities such as images and acoustics. In MiT, we also design a novel adaptive rescaling strategy at the attention head level, which optimizes the representation of infused multimodal features. Notably, all foundation models are kept frozen during the tuning process to reduce the computational burden and only 2.5% parameters are tunable. We conduct experiments across a range of multimodal tasks, including image-related tasks like referring segmentation and non-image tasks such as sentiment analysis. Our results showcase that MiT achieves state-of-the-art performance in multimodal understanding while significantly reducing computational overhead(10% of previous methods). Moreover, our tuned model exhibits robust reasoning abilities even in complex scenarios.
Figure 1: The comparison between previous multimodal tuning methods(left[23] and our proposed multimodal infusion tuning(right). The referring segmentation task is taken as an example. Previous methods usually treat visual embeddings as tokens, which are prefixed to textual tokens and some PEFT adapters are employed to tune the LLMs on downstream tasks. In our proposed multimodal infusion tuning, we infuse the visual information into each textual token in a linear manner, which is more fine-grained and dose not introduce extra tokens to pretrained LLMs.
@@ 1 Introduction
Large language models (LLMs), benefiting from their immense scale in model sizes and training data, have demonstrated impressive capabilities across a wide range of real-world tasks. The prominence of LLMs is evident in the widespread attention they have garnered from both academia and industry, with notable examples being OPT [37] and LLaMA [25]. However, training a new large model is highly resource-intensive, leading to the development of numerous parameter-efficient fine-tuning (PEFT) methods, such as LLaMA-adapter [36] and (IA)3 [19].
@@
Currently, the majority of PEFT approaches are tailored for text-only data [15, 19], posing a challenge in handling multimodal information. Empowering LLMs with multimodal data, such as images and acoustics, remains a significant challenge. To this end, some researchers have proposed integrating image information into LLMs through PEFT, exemplified by approaches like FROMAGe [12] and mPLUG-Owl [31]. Most of these methods tend to concatenate or prefix the vision embedding with text tokens and feed them directly to the LLMs, as shown in Figure 1(left). However, we identify two drawbacks in this approach. Firstly, the concatenation of tokens significantly increases memory consumption, growing quadratically due to the self-attention mechanism’s $L^{2}$ memory requirement (where $L$ represents the length of tokens). Secondly, direct prefix approaches are deemed coarse-grained, leading to insufficient interactions among representations from multiple modalities.
@@
To address these issues, we propose a new tuning framework termed multimodal infusion tuning (MiT). The whole tuning procedure is designed to be linear, ensuring low memory consumption. This infusion strategy allows us to endow LLMs with the capability to process multimodal information. One major difference between MiT and other methods is that we decouple self-attention in-depth, and inject multimodal information at a more fine-grained aspect(shown in Figure 1(right)). Specifically, we decouple the self-attention of large language models and progressively infuse the global representations of other modalities into the keys and values of text embedding. Additionally, we introduce a novel adaptive rescaling strategy at the attention head level to enhance collaboration between different modalities. We freeze the parameters in pretrained models and exclusively fine-tune those newly introduced in MiT(around only 2.5% of all parameters). Another distinguishing feature of our approach compared to previous methods is its versatility. MiT is not limited to processing image data; it can also effectively handle other types of data, including acoustic and facial features. For various multimodal tasks, we employ different downstream networks during tuning, such as a lightweight decoder for referential segmentation and a fully connected layer for sentiment analysis.
We conduct experiments on seven datasets encompassing three tasks(referring segmentation, image-text classification, and sentiment analysis), and the results demonstrate that MiT achieves state-of-the-art performance while maintaining efficiency(our proposed method requires only 0.47 TFLOPs). Moreover, leveraging the powerful language modeling capabilities of LLMs, our MiT is capable of performing reasoning and segmentation tasks guided by complex texts. For instance, the model can correctly segment phrases like spicy vegetables instead of just peppers, or best friend of human instead of simply dog.
In summary, our paper presents four key contributions:
@@
  • We introduce a fine-grained multimodal tuning strategy for LLMs, named MiT. This strategy is designed to efficiently accommodate various types of multimodal data, allowing for their progressive infusion into LLMs.
@@
  • Our tuned model demonstrates sophisticated reasoning capabilities in complex language scenarios. This suggests that the MiT approach can be effectively applied to a wider range of complex tasks and scenarios, representing the effectiveness of our method.
@@ 2 Related Works
Our work builds upon existing LLM and multimodal fine-tuning methods. In this section, we will introduce some previous methods relevant to our work.
2.1 Large Language Models and Tuning
In recent years, large language models (LLMs) have demonstrated significant advancements in in-context learning and long-term generation, as showcased by models such as ChatGPT, OPT [37], and LLaMA [25]. These models have significantly improved the capabilities of language understanding and generation. To further enhance the application of LLMs in downstream tasks, various methods have been introduced. For example, adapter-tuning [7] and prefix-tuning  [15] are two prominent techniques that focus on fine-tuning the model while preserving the original parameters. This approach offers two key benefits: it reduces computational overhead and maintains the proficiency of the LLM trained on large corpora.
Recently, there has been a surge in novel Prompt Encoding for Fine-Tuning (PEFT) methods, such as LLaMA-adapter [36] and (IA)3 [19]. Notably, (IA)3 introduces learnable vectors that are multiplied with the keys (K) and values (V) in self-attention mechanisms, effectively reducing memory usage. Our proposed method draws considerable inspiration from (IA)3, although it is specifically tailored for multimodal data and tasks. This adaptation allows our method to leverage the strengths of (IA)3 while addressing the unique challenges posed by multimodal learning.
Figure 2: The pipeline of our proposed method(left) and the detailed structure of MiT(right). The referring segmentation is shown as the example, which is part of a broader framework that includes several other tasks, such as image-text classification and sentiment analysis. In MiT, we infuse the visual information in both self-attention and feedforward module. The procedure is kept linear for computational burden consideration.
@@ 3 Multimodal Infusion Tuning
Our method seamlessly integrates multimodal information into a pretrained LLM while keeping its parameters frozen. The overall pipeline is illustrated in Figure 2(left). We introduce a tunable MiT module into the pretrained LLM, facilitating the integration of representations from other modalities. Through tuning, the LLM progressively acquires the capability to process multimodal signals as MiT module is zero-initialized. The resulting interacted textual features, combined with visual features in tasks involving images, are then fed into various downstream networks for predictions. Consequently, our method finds applications across diverse domains, including referring segmentation, image-text classification, and sentiment analysis(Figure 2 shows the example of referring segmentation).
@@ 3.1 Architecture Design
As illustrated in Figure 2(right), we infuse the multimodal information in both self-attention and feed-forward module of pretrained LLMs. We design MiT as a plugable module, so we can leverage the capabilities of the LLM learnt from large scale corpus.
Without loss of generality, assume that we have an global image representation extracted by CLIP image encoder [22]: $I\in R^{d_{I}}$, where $d_{I}$ is the corresponding feature dimension. In textual self-attention of LLM, the text embeddings are projected into query($Q_{t}$), key($K_{t}$), and value($V_{t}$). We delve into textual self-attention and integrate $I$ into $K_{t}$ and $V_{t}$. Take $V_{t}$ for example, to fit it into the space of text representations, we employ two transformations:
@@
where $W^{v}_{d},W^{v}_{a}\in R^{d_{I}\times d_{T}}$, $b^{v}_{d},b^{v}_{a}\in R^{d_{T}}$ are learnable matrices and vectors. In order to keep the computational burden linear, we infuse the transformed representation into $V_{t}$ via element-wise multiplication and addition(the InfusionV in Figure 2):
@@
where $I_{d}^{k}$ and $I_{a}^{k}$ are generated in the same approach as Equation LABEL:equ:image_trans(InfusionK in Figure 2).
@@ 3.2 Downstream Task Tuning
We fine-tune our MiT on three tasks for multimodal understanding: referring segmentation, image-text classification, and multimodal sentiment analysis. Among these tasks, sentiment analysis does not involve images, as it focuses on analyzing text or other forms of non-image data (for privacy considerations, only processed facial features are provided in the datasets evaluated). In contrast, the other two tasks, object detection and image classification, require image data for analysis. For tasks involving images, we utilize a frozen CLIP encoder to generate global semantic representations and features of images at different levels, as illustrated in Figure 2. Alternatively, for tasks not involving images, we employ a trainable transformer feature encoder to extract multimodal embeddings. Since the LLMs employed in our approach are decoder-only structures, we extract the semantic representations for downstream tasks from the last token of the entire sequences.
For referring segmentation, the goal is to segment the object from the image according to a descriptive. Under this scenario, we design a light-weight decoder, which takes infused text embedding and multi-level image features as input, and generate the predicted mask. The DICE loss [21] is employed in our framework:
@@ 3.3 Implementation Details
We evaluate our method on RefCOCO [32], RefCOCO+ [32], and G-Ref [20] for referring segmentation. There are 19994, 19992, and 26711 images, with 50000, 49856, and 54822 references, respectively. For image-text classification, we employ two datasets: UPMC-Food101 [27] and SNLI-VE [29], which contain 90840 and 565286 image-text pairs, respectively. UPMC-Food101 provides 101-category classification task while SNLI-VE aims to reason the semantic relationship(entailment, neutral, or contradiction) between text hypothesis and image premise. As for the sentiment analysis, we use two popular dataset: MOSI [33] and MOSEI [35], in which the sentiment tendency are annotated within [-3, +3]. The goal is to predict the sentiment tendency based on corresponding texts, acoustic and facial features.
We use the popular LLaMA [25] with 7B parameters as the textual foundation model. For image branch, we employ the CLIP ViT-L-336 model [22] to generate the global representations. In MiT, we infuse the multimodal representation into last certain layers in LLM, and the detailed effect are discussed in Section 5.1. There are three layers in the light-weight decoder for referring segmentation to maintain efficiency. In sentiment analysis task, three-layer transformer blocks are employed to encoder acoustic and facial features. Inherited from pretrained models, $d_{I}$, $d_{T}$, and $d_{T^{\prime}}$ are set to 768, 4096, and 11008, respectively.
@@ 4.1 Results on Referring Segmentation
Table 2: The results of our method on image-text classification benchmarks. The accuracy is utilized as metrics in the table.
@@ 5.1 Ablation Studies
Modules of MiT. There are several components in our proposed MiT, including K/V infusion(K/V-Inf.), adaptive rescaling(A.R.), feed-forward infusion(FF-Inf.), etc. The ablation study of each part on RefCOCO benchmark are shown in Table 4. As we can see from the results, K/V infusion plays the most important role in MiT. When the adaptive head rescaling is removed, the performance of the model will also decrease accordingly, demonstrating the effect of this part. Besides, infusing in feed-forward module also plays a positive role for the final results.